waths the solution to 5x -4x + 3=9 - x
Found 2 solutions by CertifiedGenius, superkid208:
Answer by CertifiedGenius(6)

I am assuming we are solving for x
5x - 4x + 3 = 9 - x [given]
x + 3 = 9 - x [subtract 4x from 5x to make 1x]
2x + 3 = 9 [add an x to both sides to collect like terms]
2x = 6 [subtract 3 from each side to collect like terms]
x = 3 [divide by 2 on each side to solve for x]

Answer by superkid208(3)

5x-4x+3=9-x
first I would subtract 5x-4x giving me x+3=9-x then I would add x to both sides of the problem giving me x+x+3=9-x+x which becomes 2x+3=9 then I would subtract 3 from both sides which becomes 2x+3-3=9-3 or 2x=6 then I would write the problem as follows 2/2x = 6/2 solving it x=3